# Advances in Tabulating Carmichael Numbers

### v1

@@ 4.4. Benefits and Crossover
It also makes deciding when to use these algorithms a bit easier. There are two upper-bounds on $q$ given $P$ and $B$. One the one hand $q<2P^{2}$ and on the other hand $Pq^{2}<B$. The cost of finding $q$ is roughly linear in the upper-bound. We can find $q$ (and $r$) by checking, on average, $O(P(\log P)^{2})$ candidate pairs $D\Delta$. Or we can find $q$ by exhaustive search on a precomputed lists of primes bounded by $\sqrt{B/P}$. The number of elements in this list is bounded by
$\pi(\sqrt{B/P})=\frac{\sqrt{B/P}}{\log\sqrt{B/P}}.$
These two counts are equalized, roughly, at $P=B^{1/3}/(\log B)^{2}$. In practice, we used an even simpler guide and treated $P<B^{1/3}$ as an unconditional bound for which we accounted as small. Note that this comparison for treating $P$ as small only takes into consideration the cost of finding $q$. Given $P$ and $q$, we still have not accounted for the cost of finding $r$. We will use, as a guide, that the cost of finding $r$ given $Pq$ is a polynomial time algorithm. In so far as our guide is wrong, it only justifies cross-over beyond $B^{1/3}$.
@@ 5. When $P$ is large
We show that the cost of completing individual preproducts $Pq$ may be exponential (see Theorem 6) but that many of these may be resolved in polynomial-time (see subsections 5.3 and 5.4 ). Even the cases that may not be resolved in polynomial-time have asymptotically superior algorithms (see Theorem 7) compared to Pinch’s approach. The difficulty of finding $r$ given $P$ and $q$ is related to both $B$ and $\lambda(Pq)$. In all of what follows, the goal is to use the two conditions on $r-1$:
@@
In what follows, we show what the bounded case is the easy case. We follow up by explaining how Pinch offered a balanced approach for preproducts not in the easy case. We provide an asymptotic justification for why this is an exponential time algorithm (see Theorem 6). The next two subsections establish several improvements each of which would change many of these individual exponential time computations to polynomial time ones. Finally, we conclude with some comments on how Carmichael numbers that result from Erdős construction technique result in the worst case inputs for this tabulation problem. A tabulation algorithm should be able to find such numbers and we find it especially pleasing that they play a central role (as an obstacle!) in a tabulation algorithm.
@@ 5.1. Pinch’s Easy Case
Using the congruence on $r-1$, we may takes steps in an arithmetic progression to find candidate $r$ values. The number of steps required is bounded by
$\min\left\{\frac{Pq-1}{\lambda(Pq)},\frac{B}{Pq\lambda(Pq)}\right\}.$
The first term comes from the fact that $r$ is a factor of $Pq-1$ and the second term comes from the fact that $r<B/Pq$. If either of these numbers is small, we just take the steps in arithmetic progression and are done. This approach may be regarded as the bounded approach. One of the two relevant bounds has forced the problem to be “easy.” For the rest of the section, we will assume that this approach is not taken. Since the bounds are not in effect, we treat the remainder as unbounded.
@@ 5.2. Pinch and Trial Division
may be sped up considerably by running over small factors $f$ of $Pq-1$. Put $r=(Pq-1)/f+1$, then test the congruence. Testing small factors $f$ lowers the upper bound on the arithmetic progression.
@@
$O\left(k+\frac{Pq-1}{k\lambda(Pq)}\right),$
where the first term represents the work for trial division up to $k$, and the second term counts the number of elements in arithmetic progression with common difference of $\lambda{(Pq)}$ that are bounded in size by $(Pq-1)/k$. These two terms are balanced at
@@ 5.3. Divisors in Residue classes (Part 1)
The search for divisors $d$ of $\mathcal{P}$ may be broken into two cases. Case 1: $d\leq\sqrt{\mathcal{P}}$. Case 2: $d>\sqrt{\mathcal{P}}$. In naive trial division, any successful division always yields two factors: one in each of the two cases. Since we are only concerned with divisors in specific residue classes, the two cases are treated independently.
@@ 5.3.1. Case 1
This does not deviate from Pinch’s approach and from the point of view of implementation, the renaming and re-scaling is not necessary for this case.
@@ 5.3.2. Case 2
Suppose we have $d>\sqrt{\mathcal{P}}$ and $d\equiv\mathcal{R}_{1}\pmod{\mathcal{S}}$. That means that there is a factor, say $d^{\prime}$ with $d(d^{\prime})=\mathcal{P}$ and $d^{\prime}<\sqrt{\mathcal{P}}$. This $d^{\prime}$ lies in some residue class modulo $\mathcal{S}$. If $(\mathcal{R}_{1}+k_{1}\mathcal{S})\mid\mathcal{P}$ then there exists $d^{\prime}=\mathcal{R}_{2}+k_{2}\mathcal{S}$ such that $(\mathcal{R}_{1}+k_{1}\mathcal{S})(\mathcal{R}_{2}+k_{2}\mathcal{S})=\mathcal{P}$. A simple computation shows that $\mathcal{R}_{2}\equiv\mathcal{P}\mathcal{R}_{1}^{-1}\pmod{\mathcal{S}}$. We construct divisors in the arithmetic progression
@@
while these candidates remain less than $\sqrt{\mathcal{P}}$. At each point, we check if $\mathcal{P}\equiv 0\mod{(\mathcal{R}_{2}+k_{2}\mathcal{S})}$. If it is, then $d=\mathcal{P}/(\mathcal{R}_{2}+k_{2}\mathcal{S})$ is the divisor greater than $\sqrt{\mathcal{P}}$. Multiply the quantity by $g$ and add $1$ to get $r$. This the same as searching for $r$ in
@@
This incorporates an insight of [9, 2] that if $\mathcal{S}>\mathcal{P}^{1/2}$, then there are only two possible divisors (those being $\mathcal{R}_{1}$ and $\mathcal{P}/\mathcal{R}_{2}$) and that they are found in polynomial time.
@@ 5.4. Divisors in Residue classes (Part 2)
If $\mathcal{S}>\mathcal{P}^{1/3}$, then there is also a polynomial-time algorithm to find all possible factors that are $\mathcal{R}_{1}\pmod{\mathcal{S}}$ based off of solving a logarithmic number of quadratic equations [9]. This result was improved in [2]; they used LLL lattice reduction techniques to find factors in polynomial time if $\mathcal{S}>\mathcal{P}^{1/4}$. In [3], the authors of [2] are reported to have said that it is practical for $\mathcal{S}>\mathcal{P}^{0.29}$.
@@ 5.5. Worst Case Inputs
Suppose $Pq$ is a Carmichael number. In this case, $r^{\star}-1=0$ and so we have no “residue class” information to use; the problem is to factor $(Pq-1)/\lambda(Pq)$. By the Erdős construct techniques, it is possible to make Carmichael numbers with $\lambda(Pq)$ arbitrarily small as a power of $Pq$. To make matters worse, $Pq$ may be a Carmichael number less than $B^{2/3}$. In this case, there is no relevant bound information. It is truly a general integer factorization problem. Since this is a worst case input to our algorithm, it would make sense to analyse the asymptotic density of these numbers to know how often we encounter them. On the one hand, we could consult current tabulations as a guide for how often we expect to have these inputs. On the other hand, the asymptotic count remains a key open question in analytic number theory.
@@ 6. Bound Admissible Preproducts
In sections 4 and 5, we make the case for a bifurcated approach for tabulation. In each case, we studied the problem as the following: Given $P$ find $q$ and $r$ so that $Pqr$ is a Carmicheal number. In section 4, $q$ and $r$ are found simultaneously. In section 5, we see that the tabulation bound $B$ implies that an exhaustive search for $q$, followed by a constructive search for $r$ is reasonable. In both cases, this involves finding $P$ efficiently. A bound admissible preproduct $P_{k}$ for $k<d$ is a cyclic number where $P_{k}p_{k}^{d-k}<B$. In section 7, we show that due to the count of bound-admissible preproducts, care must be taken in generating these numbers. As with the tabulation itself, our approach is bifurcated.
@@ 6.2. Bound admissible $P_{d-1}$
In section 5, we saw that $q$ is found in the interval $(p_{d-2},\sqrt{B/P})$. To account for any prime, the smallest $P$ will be is at the cross-over $X$. If $X$ is chosen to be at least $B^{1/3}$, that means the table will only require primes up to $B^{1/3}$. Naturally, the larger we choose $X$, the smaller the table needs to be.
At our scale, a significant portion of the total tabulation time of the “large” case was spent just in preproduct creation. It was important to both minimize duplicated work and yet have a balanced, efficient parallel algorithm. To properly distribute the computation of preproducts (and thus the work to find the completions of such preproducts) with $k$ prime factors, every processor created all bound-admissible preproducts with $\lfloor k/2\rfloor$ prime factors. These were counted and then a given processor advanced to the next appending more primes only when the count of preproducts is in its residue class modulo the total number of processors.
@@ 7. Asymptotic counts of Preproducts
We were initially concerned with the count of preproducts from the “small” case. We posed that problem in a session at the West Coast Number Theory Conference (2022) and Sunjum Kim showed us how to count that set. In particular, he proved the following.
@@
We believe, as a guide, that the cost of completing a given preproduct $Pq$ is a polynomial-time algorithm, and that the bound in Theorem 8 serves as a guide for the count of preproducts constructed.
A better guide might be to know the distribution of preproducts in $[0,B]$. In Section 5, we saw that the larger a preproduct was, the more likely it was to be an “easy case.” Both the number of difficult problems and their distribution is important to establish sharper asymptotic bounds.
@@ 8. Implementation, statistics, and conclusion
The algorithm in Section 3 was implemented in C++ [17]. Production code was run on Big Dawg, a cluster at Butler University. Timings were performed on Hyperion, a 48-core (3.8 GHz), cluster housed at Illinois Wesleyan University.
Code for the small preproduct case was implemented in time for the publication of [18], but tabulation was only performed on prime preproducts. We can now report that unbounded, small preproduct tabulation has been performed for preproducts up to 70 million, and bounded, large preproduct tabulation was performed up to $10^{22}$. The latter was broken into cases by $d$, the number of prime factors of the resulting Carmichael numbers, and was performed for $d=4$ through $d=13$. Taken together, this computation represents a new record in Carmichael tabulation. The total wall time of the computation was almost two years.
The following table gives timings in seconds for the pinch method and the new method.
$\begin{array}[]{|c|c|c|c|c|c|c|}\hline\cr B&d&\mbox{pinch}&\mbox{new}&d&\mbox{pinch}&\mbox{new}\\ \hline\cr 10^{14}&4&24&33&6&25&35\\ \hline\cr 10^{15}&4&125&170&6&189&242\\ \hline\cr 10^{16}&4&622&842&6&1306&1671\\ \hline\cr 10^{17}&4&3338&4347&6&8865&11627\\ \hline\cr\end{array}$
Unfortunately the new method is slower in this range. The key problem is that within these bounds, the proportion of cases where the new method gives an exponential speedup are relatively rare. Meanwhile, the polynomial time cost for new calculations are applied to every preproduct. We know that the new method outperforms Pinch’s at asymptotic scales, but unfortunately the scales at which computations are performed are not asymptotic. One observation that gives hope is that the ratio of timings (new to pinch) decreases as the bound increases.
As an example of a particular preproduct where the new method gains an advantage, consider $P=11756813235$. The timings measured were $9$ microseconds for the new method, and $17$ for Pinch’s method.

### v2

@@ 4.4. Benefits and Crossover
It also makes deciding when to use these algorithms a bit easier. There are two upper bounds on $q$ given $P$ and $B$. On the one hand $q<2P^{2}$ and on the other hand $Pq^{2}<B$. The cost of finding $q$ is roughly linear in the upper bound. We can find $q$ (and $r$) by checking, on average, $O(P(\log P)^{2})$ candidate pairs $D\Delta$. Or we can find $q$ by exhaustive search on a precomputed lists of primes bounded by $\sqrt{B/P}$. The number of elements in this list is bounded by
$\pi(\sqrt{B/P})\sim\frac{\sqrt{B/P}}{\log\sqrt{B/P}}.$
These two counts are equalized, roughly, at $P\sim B^{1/3}/(\log B)^{2}$. In practice, we recommend a crossover of $X=B^{1/3}$. Our own computation went a bit beyond this. This was due, in part, because when we started the small case, we were unsure what the upper bound of the total tabulation was going to be. As a guide, the cost of finding $r$ given $Pq$ is a polynomial time algorithm. However, if the guidance is incorrect then choosing $X$ larger than $B^{1/3}$ would be justified.
@@ 5. When $P$ is large
We show that the cost of completing individual preproducts $Pq$ may be exponential (see Theorem 6) in the worse case, but that many of these may be resolved in polynomial-time (see Subsections 5.3 and 5.4 ). Even the cases that may not be resolved in polynomial-time have asymptotically superior algorithms (see Theorem 7) compared to Pinch’s approach. The difficulty of finding $r$ given $P$ and $q$ is related to both $B$ and $\lambda(Pq)$. In all of what follows, the goal is to use the two conditions on $r-1$:
@@
The first two subsections explain the simplest approaches to finding $r$; both of which are described in [15]. We prove that this can be an exponential time algorithm in $P$ (see Theorem 6). The next two subsections establish several improvements each of which would change many of these individual exponential time computations to polynomial time ones. Finally, we conclude with some comments on how Carmichael numbers that result from the Erdős construction technique result in the worst case inputs for this tabulation problem. Our algorithm can find such numbers, but we find it especially pleasing that they play a central role (as an obstacle!) in the tabulation algorithm.
@@ 5.1. Pinch’s Easy Case
Using the congruence on $r-1$, we may take steps in an arithmetic progression to find candidate $r$ values. The number of steps required is bounded by
$k=\min\left\{\frac{Pq-1}{\lambda(Pq)},\frac{B}{Pq\lambda(Pq)}\right\}.$
The first term comes from the fact that $r$ is a factor of $Pq-1$ and the second term comes from the fact that $r<B/Pq$. If either of these numbers is small, we take the $k$ steps in this arithmetic progression. Recall our guidance that $\lambda(Pq)$ is close in size to $Pq$. Therefore, it is common for the first ratio to be small. In a similar way, when $Pq>B^{1/2}$, then the second ratio is expected to be small. Most of the preproducts fall into one of these two cases and we consider this the bounded case. For the rest of the section we assume both of these ratios are large. Therefore, everything that follows is considered to be in the unbounded case.
@@ 5.2. Pinch and Trial Division
may be sped up considerably by running over small factors $f$ of $Pq-1$ to find candidates for $r$. Put $r=(Pq-1)/f+1$, then test the congruence. Testing small factors $f$ lowers the upper bound on the arithmetic progression.
@@
$O\left(k+\frac{Pq-1}{k\lambda(Pq)}\right).$
The first term of the sum represents the number of candidates by taking $k$ steps of trial division. The second term counts the number of elements in arithmetic progression with common difference of $\lambda{(Pq)}$ that are bounded in size by $(Pq-1)/k$. These two terms are balanced at
@@ 5.3. Divisors in Residue classes (Part 1)
The search for divisors $t$ of $\mathcal{P}$ may be broken into two cases. Case 1: $t\leq\sqrt{\mathcal{P}}$. Case 2: $t>\sqrt{\mathcal{P}}$. In naive trial division, any successful division always yields two factors: one in each of the two cases. Since we are only concerned with divisors in specific residue classes, the two cases are treated independently.
@@ 5.3.1. Case 1
This arithmetic progression does not deviate from the original approach. The renaming is only necessary to compute the new upper bound.
@@ 5.3.2. Case 2
Suppose we have $t>\sqrt{\mathcal{P}}$ and $t\equiv\mathcal{R}_{1}\pmod{\mathcal{S}}$. That means that there is a factor, say $t^{\prime}$ with $t(t^{\prime})=\mathcal{P}$ and $t^{\prime}<\sqrt{\mathcal{P}}$. This $t^{\prime}$ lies in some residue class modulo $\mathcal{S}$. If $(\mathcal{R}_{1}+k_{1}\mathcal{S})\mid\mathcal{P}$ then there exists $t^{\prime}=\mathcal{R}_{2}+k_{2}\mathcal{S}$ such that $(\mathcal{R}_{1}+k_{1}\mathcal{S})(\mathcal{R}_{2}+k_{2}\mathcal{S})=\mathcal{P}$. A simple computation shows that $\mathcal{R}_{2}\equiv\mathcal{P}\mathcal{R}_{1}^{-1}\pmod{\mathcal{S}}$. We construct divisors in the arithmetic progression
@@
while these candidates remain less than $\sqrt{\mathcal{P}}$. At each point, we check if $\mathcal{P}\equiv 0\mod{(\mathcal{R}_{2}+k_{2}\mathcal{S})}$. If it is, then $t=\mathcal{P}/(\mathcal{R}_{2}+k_{2}\mathcal{S})$ is the divisor greater than $\sqrt{\mathcal{P}}$. Multiply the quantity by $g$ and add $1$ to get $r$. This is the same as searching for $r$ in
@@
This incorporates an insight of [14, 3] that if $\mathcal{S}>\mathcal{P}^{1/2}$, then there are only two possible divisors (those being $\mathcal{R}_{1}$ and $\mathcal{P}/\mathcal{R}_{2}$) and that they are found in polynomial time. Given our guidance about the size of $\lambda(Pq)$, the expected case is for there to be only two possible divisors.
@@ 5.4. Divisors in Residue classes (Part 2)
If $\mathcal{S}>\mathcal{P}^{1/3}$, then there is also a polynomial time algorithm to find all possible factors that are $\mathcal{R}_{1}\pmod{\mathcal{S}}$ based off of solving a logarithmic number of quadratic equations [14]. This result was improved in [3]; they used LLL lattice reduction techniques to find factors in polynomial time if $\mathcal{S}>\mathcal{P}^{1/4}$. In [4], the authors of [3] are reported to have said that it is practical for $\mathcal{S}>\mathcal{P}^{0.29}$.
@@ 5.5. Worst Case Inputs
Suppose $Pq$ is a Carmichael number. In this case, $r^{\star}-1=0$ and so we have no residue class information to use; the problem is to factor $(Pq-1)/\lambda(Pq)$. By the Erdős construction, it is possible to make Carmichael numbers with $\lambda(Pq)$ arbitrarily small as a power of $Pq$. To make matters worse, $Pq$ may be a Carmichael number less than $B^{2/3}$. In this case, there is no relevant bound information. It is truly a general integer factorization problem. Since this is a worst case input to our algorithm, it would make sense to analyse the asymptotic density of these numbers to know how often we encounter them. On the one hand, we could consult current tabulations as a guide for how often we expect to have these inputs. On the other hand, the asymptotic count remains a key open question in analytic number theory.
@@ 6. Bound Admissible Preproducts
In Sections 4 and 5, we make the case for a bifurcated approach for tabulation. In each case, we studied the problem as the following: Given $P$ find $q$ and $r$ so that $Pqr$ is a Carmichael number. In Section 4, $q$ and $r$ are found simultaneously. In Section 5, we see that the tabulation bound $B$ implies that an exhaustive search for $q$, followed by a constructive search for $r$ is reasonable. In both cases, this involves finding $P$ efficiently. A bound admissible preproduct $P_{k}$ for $k<d$ is a cyclic number where $P_{k}p_{k}^{d-k}<B$. In Section 7, we show that due to the count of bound-admissible preproducts, care must be taken in generating these numbers. As with the tabulation itself, our approach is bifurcated.
@@ 6.2. Bound admissible $P_{d-1}$
In Section 5, we showed that $q$ is in the interval $(p_{d-2},\sqrt{B/P})\subset(p_{d-2},\sqrt{B/X})$. If $X$ is chosen to be at least $B^{1/3}$, that means the table will only require primes up to $B^{1/3}$. Naturally, the larger we choose $X$, the smaller the table needs to be.
At our scale, a significant portion of the total tabulation time of the large case was spent just in preproduct creation. It was important to both minimize duplicated work and yet have a balanced, efficient parallel algorithm. To properly distribute the computation of preproducts (and thus the work to find the completions of such preproducts) with $k$ prime factors, every processor created all bound-admissible preproducts with $\lfloor k/2\rfloor$ prime factors. These were counted and then a given processor advanced to the next appending more primes only when the count of preproducts is in its residue class modulo the total number of processors.
@@ 7. Asymptotic counts of Preproducts
We were initially concerned with the count of preproducts from the small case. We posed that problem in a session at the West Coast Number Theory Conference (2022) and Sungjin Kim showed us how to count that set. In particular, he proved the following.
@@
We believe, as a guide, that the cost of completing a given preproduct $Pq$ is a polynomial-time algorithm, and that the bound in Theorem 8 serves as a guide for the count of preproducts constructed. A better guide might tell us the distribution of preproducts in $[0,B]$. In Section 5, we saw that the larger a preproduct was, the more likely it was to be an “easy case.” So, both the number of difficult problems and their distribution is important to establish sharper asymptotic bounds.
@@ 8. Implementation, statistics, and conclusion
While our high level overview in Section 3 presented tabulation as a single algorithm, the implementation consisted of many different programs. We chose $X=7\cdot 10^{7}$ as our cross-over point and $B=10^{22}$. The choice of $X$ was made before we knew what would be a feasible choice for $B$. The small and large preproduct cases were implemented and run independently. The large case was written so that it finds Carmichael numbers with exactly $d$ prime factors and had many separate computations: one for each $4\leq d\leq 9$ and one recursive version that handled $d>9$. We implemented in C++; everything was done in machine word size except three quantities that used GMP. In the small case, we constructed Carmichael numbers outside the machine word bounds and also used GMP for primality testing of $q$ and $r$. The code may be found on GitHub [23].
Combining large and small cases provides the exact count by prime factors for the total tabulation up to $B=10^{22}$.
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr d&3&4&5&6&7\\ \hline\cr C_{d}(10^{22})&420658&2327421&1370257&6034046&14056372\\ \hline\cr\hline\cr 8&9&10&11&12&13\\ \hline\cr 16005645&8939434&2347828&262818&10018&55\\ \hline\cr\end{array}$
Each of these numbers was proven to be a Carmichael number with a separate check program, and we confirmed that the numbers up to $10^{21}$ matched Pinch’s most recent tabulation.
The next two subsections give greater details about the two different computational regimes.
8.1. Small Preproduct Information
In [24], the computation was meant to be a proof of concept and had a relatively small bound of $P<3\cdot 10^{6}$. We have extended the bound to $P<7\cdot 10^{7}$. In Theorem 5 of [24], it was shown that the total cost to use all $P<X$ is $O((X\log X)^{2})$. The small case tabulation here represents about $800$ times more work over that of [24]. It was computed using 33.1 vCPU years on AMD EPYC 7B13 CPUs [^1] Special thanks to Drew Sutherland and the Simons Collaboration in Arithmetic Geometry, Number Theory, and Computation for computational support..
We did the computation unbounded and found $35985331$ Carmichael numbers. Of these, only $1202914$ were less than $B$. In the table below, we show the contribution to the tabulation by order of magnitude:
$\begin{array}[]{|c|c|c|c|c|c|c|}\hline\cr 10^{3}&10^{4}&10^{5}&10^{6}&10^{7}&10^{8}&10^{9}\\ \hline\cr 1&7&16&43&105&255&646\\ \hline\cr\hline\cr 10^{10}&10^{11}&10^{12}&10^{13}&10^{14}&10^{15}&10^{16}\\ \hline\cr 1547&3605&8238&19019&40772&79417&137794\\ \hline\cr\hline\cr 10^{17}&10^{18}&10^{19}&10^{20}&10^{21}&10^{22}&10^{23}\\ \hline\cr 219887&327732&464605&639763&872861&1202914&1708687\\ \hline\cr\end{array}$
It should be noted that this produced a complete tabulation of all Carmichael numbers less than $10^{11}$. For the $10^{12}$ bound, this computation only missed three Carmichael numbers:
  • $702712420201=(11\cdot 37\cdot 43\cdot 61\cdot 71)\cdot 73\cdot 127$,
  • $919707221161=(19\cdot 29\cdot 43\cdot 53\cdot 73)\cdot 79\cdot 127$,
  • $995483689201=(19\cdot 31\cdot 37\cdot 53\cdot 61)\cdot 71\cdot 199$.
Each of these numbers has a preproduct (the portion in parentheses) that exceeds $7\cdot 10^{7}$. So, this tabulation regime could not have found these numbers, but they were found in the large preproduct portion of the tabulation with $d=7$. For the $10^{13}$ bound, there are $260$ missing Carmichael numbers. And so it goes; the small case made the most significant contributions with smaller $d$ values. For the cases $d\geq 7$, the small case found no Carmichael numbers.
Since we ran this part of the computation as unbounded, we found many Carmichael numbers that were significantly larger than $10^{22}$. The largest one we found is
699991334899878690750821171493630078866294519097=699991334899878690750821171493630078866294519097 69999133$\cdot$ 4899878690750821$\cdot$ 171493630078866294519097=69999133 ⋅ 4899878690750821 ⋅ 171493630078866294519097 =
58 82013 03152 54068 53935 58087 37155 82013 87008 71721.58820130315254068539355808737155820138700871721 58\ 82013\ 03152\ 54068\ 53935\ 58087\ 37155\ 82013\ 87008\ 71721.58 82013 03152 54068 53935 58087 37155 82013 87008 71721 .
This example may be reproduced with the parameters $P=69999133,D=2,$ and $\Delta=1$ and would have been found with the $D\Delta$ method. The dynamic choice in the hybrid method would have chosen the lesser amount of work: iterate through $768$ divisors or iterate through approximately $7\cdot 10^{7}$ values of $C$. This example represents $q$ and $r$ taking maximal values. That is, $q-1=(P-1)(P+2)$ and $r-1=(Pq-1)/2$.
8.2. Large Preproduct Information